# Quasi Instrumental Variable Methods for Stable Hidden Confounding and Binary Outcomes

### v1

@@ 3.3 Semiparametric Locally Efficient Triply Robust Estimation
As in practice, one may not be certain about the posited baseline outcome model $p_{00}(x,z)$, we consider an alternative potentially more robust estimator to partially guard against model mis-specification. In this vein, we now propose an estimator that is consistent for the causal effect even if the outcome regression model $p_{00}(x,z)$ is not correctly specified, provided that working models for $\alpha({x})$ and
@@
are correctly specified. Note that modeling P(A,Z|X)P(A,Z|{X}) does not involve the outcome $Y$, and thus can be carried out separately from the MLE using only data on $(A,Z,X)$. In order to obtain the desired triply robust estimator, we derive the efficient influence function (EIF) of $\gamma$ in the nonparametric model in which no priori restriction is placed on the observed data law (Newey, 1990; Bickel et al., 1993).
@@
To establish triply robustness of the efficient influence function, consider the following three semiparametric models $\mathcal{M}_{1}$, $\mathcal{M}_{2}$, and $\mathcal{M}_{3}$:
@@
In the proof for Theorem 4, we establish that the moment condition $E\{\psi(O;\gamma)\}=0$ holds in the union model of $\mathcal{M}_{1}$, $\mathcal{M}_{2}$ and $\mathcal{M}_{3}$, and thus under standard regularity conditions the corresponding moment-based estimator for the marginal ATT $\gamma$ that solves an empirical version of the moment condition,
@@
The following Theorem 4 summarizes the properties of our proposed triply robust estimator $\widehat{\gamma}_{tr}$, and the proof is provided in the Supplementary Materials.
@@ 3.4 Many Weak Quasi-Instruments
Multiple QIVs can be incorporated under the proposed framework to improve identification and efficiency. For the maximum likelihood estimation in Section 3.2, when $m$ genetic variants are available, we can replace the scalar $Z$ in the log-likelihood (4) by a $m$-dimensional vector $\boldsymbol{Z}$. Let $\hat{\varphi}$ be the solution to the corresponding score function, i.e., $S_{n}({\varphi})=\partial\ell_{n}(\varphi)/\partial\varphi=0$; and let
@@
as the sample size $n\rightarrow+\infty$. In other words, the MLE is consistent and asymptotically normal if the total identification signal is strong enough. In practice, we can empirically evaluate this condition using the ratio between the minimum eigenvalue of the negative Hessian matrix and the number of parameters: $\hat{\kappa}=-\lambda_{\min}\{I_{n}(\hat{\varphi})\}/k$. In practice, we recommend checking that the $\hat{\kappa}$ is at least greater than 10 (Liu et al., 2023). For the semiparametric estimation in Section 3.3, a fixed number of semiparametric moments can be incorporated by adopting a standard generalized method of moments (GMM) approach (Hansen, 1982). We leave it as future work to consider the setting where the number of moments is allowed to diverge.
@@ 5 Real Data Analysis: Adiposity and Hypertension in the UK Biobank
The exposure of interest was intermediate adiposity (overweight and obese), defined as a BMI greater than 25 kg/m2 (Clarke et al., 2015; Timpson et al., 2009). The outcome was hypertension, defined as SBP $\geq$ 140 mmHg or DBP $\geq$ 90 mmHg. Among eligible individuals, the mean age was 57 years (SD = 7.98), 46.3% were male, and the average BMI was 27.37 kg/m2. Approximately 66.6% were classified as overweight, and 51.2% met criteria for hypertension.
@@
The three SNPs—rs3817334, rs3888190, and rs11126666—were selected from a previously published Mendelian randomization study of adiposity and cardiometabolic risk (Sun et al., 2019), based on their documented relevance to adiposity-related pathways. To investigate the potential impact of residual unmeasured confounding, we use these SNPs as QIVs. For each SNP, we estimated the average treatment effect on the treated (ATT) using both the maximum likelihood estimator (MLE) and the proposed triply robust (TR) estimator, adjusting for sex and age. The MLE estimates were tightly clustered across SNPs, ranging from 0.0948 to 0.0962, reflecting stability under the parametric modeling assumptions. In contrast, the TR estimates ranged from 0.0863 to 0.1286, exhibiting moderate variability across SNPs. This variation is expected given that the TR estimator integrates information from multiple nuisance models and remains consistent under broader conditions, offering protection against model misspecification. Notably, all TR estimates were lower than the fully adjusted g-formula estimate, suggesting that traditional outcome regression methods may still be subject to unmeasured confounding. Across all three SNPs, the marginal $\alpha$ estimates were consistently close to 1.14, indicating modest upward confounding—i.e., a tendency for conventional estimates to overstate the causal effect of overweight on hypertension risk. Figure 7 summarizes the point estimates and confidence intervals for the ATT estimates obtained from the g-formula and QIV methods.
Figure 7: Forest plot of ATT estimates for the effect of overweight on hypertension risk using the g-formula and QIV methods. Error bars represent 95% confidence intervals. The g-formula results (unadjusted, sex-adjusted, and fully adjusted) are presented first, followed by SNP-specific ATT estimates using the QIV maximum likelihood estimator (MLE) and triply robust (TR) estimator. All QIV analyses are adjusted for age and sex.
@@
Importantly, the TR estimates were consistently lower than the fully adjusted g-formula estimate, providing evidence that residual confounding—uncorrected even after covariate adjustment—may still inflate standard regression-based estimates. The QIV framework, by leveraging genetic variants and semiparametric estimation, provides a more credible approach to causal inference when standard adjustment methods fall short.

### v2

@@ 3.3 Semiparametric Locally Efficient Triply Robust Estimation
Often, one may not be certain that the posited baseline outcome model is correctly specified. Therefore, we consider an alternative potentially more robust estimator to partially guard against model misspecification. In this vein, we now propose an estimator that is consistent for the causal effect even if the outcome regression model $p_{00}(x,z)$ is not correctly specified, provided that working models for $\alpha({x})$ and
@@
are correctly specified. Note that modeling P(A,Z|X)P(A,Z|{X}) does not involve the outcome $Y$, and thus can be carried out separately from the MLE using only data on $(A,Z,X)$. In order to obtain the desired triply robust estimator, we derive the efficient influence function (EIF) of $\gamma$ in the nonparametric model in which no prior restriction is placed on the observed data law (Newey, 1990; Bickel et al., 1993).
@@
The proof of Theorem 3 is provided in Section S4.3 of the Supplementary Materials. To establish triply robustness of the efficient influence function, consider the following three semiparametric models $\mathcal{M}_{1}$, $\mathcal{M}_{2}$, and $\mathcal{M}_{3}$:
@@
In the proof of Theorem 4 (see Section S4.4 of the Supplementary Materials), we establish that the moment condition $E\{\psi(O;\gamma)\}=0$ holds in the union model $\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3}$. Under standard regularity conditions, this implies that the corresponding moment-based estimator for the marginal ATT $\gamma$ is consistent when solving the empirical version of the moment condition,
@@
Theorem 4 summarizes the properties of our proposed triply robust estimator $\widehat{\gamma}_{tr}$. The proof is given in Section S4.4 of the Supplementary Materials.
@@ 3.4 Many Weak Quasi-Instruments
Multiple QIVs can be incorporated under the proposed framework to improve identification and efficiency. For example, in an epidemiological analysis using genetic variants as QIV in the maximum likelihood estimation approach in Section 3.2 where $m$ genetic variants are available, one can replace the scalar $Z$ in the log-likelihood (4) by a $m$-dimensional vector $\boldsymbol{Z}$. Let $\hat{\varphi}$ be the solution to the corresponding score function, i.e., $S_{n}({\varphi})=\partial\ell_{n}(\varphi)/\partial\varphi=0$; and let
@@
as the sample size $n\rightarrow+\infty$. In other words, the MLE is consistent and asymptotically normal if the total identification signal is strong enough. In practice, we can empirically evaluate this condition using the ratio between the minimum eigenvalue of the negative Hessian matrix and the number of parameters: $\hat{\kappa}=-\lambda_{\min}\{I_{n}(\hat{\varphi})\}/k$. In practice, we recommend checking that the $\hat{\kappa}$ is at least greater than 10 (Liu et al., 2023). For the semiparametric estimation in Section 3.3, a fixed number of semiparametric moments can be incorporated by adopting a standard generalized method of moments (GMM) approach (Hansen, 1982). We leave it as future work to consider the setting where the number of moments is allowed to diverge. Further technical details on the extension to settings with many (possibly weak) QIVs are provided in Section S4.5 of the Supplementary Materials.
@@ 5 Real Data Analysis: Adiposity and Hypertension in the UK Biobank
The exposure of interest was a binary indicator of intermediate adiposity (overweight and obese), defined as a BMI greater than 25 kg/m2 (Clarke et al., 2015; Timpson et al., 2009). The outcome was hypertension, defined as SBP $\geq$ 140 mmHg or DBP $\geq$ 90 mmHg. Among eligible individuals, the mean age was 57 years (SD = 7.98), 46.3% were male, and the average BMI was 27.37 kg/m2. Approximately 66.6% were classified as overweight, and 51.2% met criteria for hypertension.
@@
The three SNPs—rs3817334, rs3888190, and rs11126666—were selected from a previously published Mendelian randomization study of adiposity and cardiometabolic risk (Sun et al., 2019), based on their documented relevance to adiposity-related pathways. To investigate the potential impact of residual unmeasured confounding, we use these SNPs as QIVs. For each SNP, we estimated the average treatment effect on the treated (ATT) using both the maximum likelihood estimator (MLE) and the proposed triply robust (TR) estimator, adjusting for sex and age. The MLE estimates were tightly clustered across SNPs, ranging from 0.0948 to 0.0962, reflecting stability under the parametric modeling assumptions. In contrast, the TR estimates ranged from 0.0863 to 0.1286, exhibiting moderate variability across SNPs. This variation is expected given that the TR estimator integrates information from multiple nuisance models and remains consistent under broader conditions, offering protection against model misspecification. Notably, all TR estimates were lower than that of the fully adjusted g-computation, suggesting that traditional outcome regression methods may still be subject to unmeasured confounding. Across all three SNPs, the marginal $\alpha$ estimates were consistently close to 1.14, indicating modest upward confounding—i.e., a tendency for conventional estimates to overstate the causal effect of overweight on hypertension risk. Figure 7 summarizes the point estimates and confidence intervals for the ATT estimates obtained from the g-computation and QIV methods.
Figure 7: Forest plot of ATT estimates for the effect of overweight on hypertension risk using the g-computation and QIV methods. Error bars represent 95% confidence intervals. The g-computation results (unadjusted, sex-adjusted, and more covariates) are presented first, followed by SNP-specific ATT estimates using the QIV maximum likelihood estimator (MLE) and triply robust (TR) estimator. All QIV analyses are adjusted for age and sex.
@@
Importantly, the TR estimates were consistently lower than that of the adjusted g-computation with more covariates, providing evidence that residual confounding, uncorrected even after covariate adjustment, may still inflate standard regression-based estimates. As shown in Figure 7, both the pooled TR (GMM) and pooled MLE estimates provide consistent evidence of a positive causal effect. For comparison, we also examined the classical two-stage least squares (2SLS) estimator using all three SNPs jointly. The 2SLS point estimate (0.173, 95% CI: [0.014,0.360][-0.014,0.360]) was close to the unadjusted g-computation estimate, suggesting that it likely did not remove unmeasured confounding. Moreover, the wide interval covering the null reflects the weakness of these instruments under the conventional IV framework. It is important to note that our three selected SNPs were chosen under the QIV framework, not the classical IV assumptions, and therefore may fail to meet IV relevance or other core IV conditions. Our proposed QIV framework, by leveraging genetic variants and semiparametric estimation, provides a credible complementary approach to causal inference in addition to standard adjustment methods or 2SLS. Additional empirical results, including two-stage least squares (2SLS) estimates, are provided in Section S5 of the Supplementary Materials.